probability
1) Suppose your chance of being offered a certain job is 0.45, your probability of getting another job is 0.55, and your probability of being offered both jobs is 0.30. What is the probability that you will be offered at least one of the two jobs?

probabiliti distribution
1) The lifespans of people in a certain city constitute a normal distribution with a mean of 72 years and a standard deviation of 6 years. What is the probability that a randomly selected person from the city will live more than 75 years
Answer #1

(1)

By Addition Theorem:
P(At least one of the two jobs) = P( being offered a certain job ) + P( getting another job is 0.55) - P( being offered both jobs )

= 0.45 + 0.55 - 0.30

= 0.70

So,

Answer is:

0.70

(1)

\mu = 72

\sigma = 6

To find P(X>75):

Z = (75 - 72)/6

= 0.50

By Technology, Cumulative Area Under Standard Normal Curve = 0.6915

So,

P(X>75) = 1 - 0.6915 = 0.3085

So,

Answer is:

0.3085